# Augmenting LLMs for General Time Series Understanding and Prediction

### v1

@@ Abstract
Time series data is fundamental to decision-making in many crucial domains including healthcare, finance, and environmental science. However, analyzing this data often requires incorporating unstructured contextual information, answering domain-specific questions, and generating natural language explanations – capabilities that traditional time series models lack due to their inability to process text. While Large Language Models (LLMs) excel at contextual reasoning and knowledge integration, they struggle with numerical time series due to inefficient text-based representations and limited exposure to temporal data during pretraining. We address this gap by augmenting an LLM with specialized time series perception through a patch-based encoder-decoder architecture. We train this Time Series-augmented LLM (TsLLM) on a large corpus of over 2 million interleaved time series and text examples spanning diverse analysis tasks: forecasting with contextual information, time series question-answering, pattern explanation, classification with natural language outputs, and report generation. This training enables TsLLM to leverage both its language understanding and newly acquired temporal reasoning capabilities. While not designed to surpass specialized models on traditional benchmarks, TsLLM demonstrates strong performance on tasks requiring the integration of time series analysis with natural language – capabilities that existing approaches cannot provide. Our work establishes a new paradigm for time series analysis that bridges numerical computation and natural language understanding, democratizing access to sophisticated temporal reasoning through natural language interaction.
@@ 1 Introduction
Many real-world time series workflows require more than numerical prediction. Clinicians ask questions about electrocardiograms in the context of symptoms and medications, analysts forecast while conditioning on news and product feedback, and scientists need narrative explanations of temporal patterns and anomalies. Traditional time series models are effective for narrowly defined numerical tasks, but they operate exclusively on structured inputs and produce only numerical outputs. As a result, practitioners often rely on brittle pipelines or expert intermediaries to bridge the gap between temporal signals, unstructured context, and high-level analysis. Furthermore, these models require data science expertise to build and interpret, presenting a significant barrier to entry for domain experts. Across healthcare, finance, scientific research, and industrial monitoring, domain experts need systems that can reason about time series data through natural language, answer questions about temporal patterns, and provide explanations for their analyses. The rise of multimodal machine learning models has transformed how we interact with visual and audio data, yet time series remains isolated from these advances, despite being one of the most pervasive data modalities in scientific and medical settings.
@@
Large language models (LLMs) possess the capabilities needed for these integrated tasks: contextual reasoning, few-shot learning from examples, domain knowledge integration, and natural language interaction. As a result, utilizing LLMs to perform certain time series analysis tasks holds considerable promise (Jin et al., 2024; Zhang et al., 2024). However, their weak abilities to perceive numerical time series data prevent LLMs from effectively applying their capabilities to temporal analysis (Tan et al., 2024; Merrill et al., 2024). Two structural issues explain this gap. First, text tokenization represents each numeric value with multiple tokens, inflating sequence length and obscuring temporal structure; this limits the amount of signal and context that fits in context windows and hinders pattern recognition. Second, our analysis of LLM pretraining corpora reveals they contain little high-quality time series data, so LLMs lack the exposure needed for temporal analysis.
@@
To address these limitations, we introduce TsLLM, a large language model augmented with dedicated time series perception that enables unified, interleaved text-time series reasoning and generation. TsLLM adds a high-fidelity, patch-based encoder-decoder pretrained with a $\beta$-VAE objective to produce compact, information-dense representations of temporal signals, along with a scale-aware encoding that decouples shape from magnitude so the model can reason about absolute scale while maintaining numerical stability. Lightweight MLP adapters map these time series latents into the LLM’s embedding space, and special tokens allow seamless fusion: inputs and outputs can interleave text and time series segments, and the model treats all tasks as next-token prediction. This preserves the pretrained LLM’s language ability while adding precise temporal perception. TsLLM’s unified formulation of language and time series modeling enables functions that traditional time series models and text-only LLMs generally do not provide: contextual forecasting conditioned on unstructured information, question answering over numerical time series data, and natural language guided time series synthesis, as well as standard time series tasks: forecasting, classification, anomaly detection and more, without using task-specific heads. Few-shot in-context prompting with interleaved examples further allows the model to adapt to new datasets and tasks without additional fine-tuning.
@@
Empirically, TsLLM demonstrates strong performance on tasks that require integrating temporal signals with language. TsLLM significantly outperforms frontier LLMs on time series question-answering (QA) benchmarks, matches or exceeds specialized systems for QA on electrocardiogram data, and beats both LLM-based methods and traditional models on a contextual forecasting benchmark. It also retains most of its performance on standard text-only LLM evaluations. On traditional time series benchmarks, TsLLM is competitive with dedicated models on forecasting (e.g., GIFT-Eval) and classification (e.g., UCR), even though it is not optimized for purely numerical tasks. Ablations show that few-shot prompting provides significant gains in most settings, and that the performance of TsLLM scales with the size of the pretrained LLM it is built on, suggesting that its performance can be improved by scaling up further. The primary contributions of this work are:
@@
    A large corpus of multimodal interleaved text and numerical time series sequences, comprising over two million examples and 20 billion tokens in total, which were collected, augmented, and synthesized for training TsLLM.
@@
    We are releasing both the data and code for TsLLM to enable further research in this area (links to come).
@@ 2 Related Works
Multimodal LLMs
Multimodal Large Language Models (MLLMs) extend LLMs’ capabilities beyond text to process diverse data types (Yin et al., 2024). These models integrate modality-specific encoders (e.g., vision encoders like CLIP (Radford et al., 2021) or audio encoders like CLAP (Elizalde et al., 2023)), with LLM backbones (e.g. LLaMa (Touvron et al., 2023; Grattafiori et al., 2024) and Qwen (Wang et al., 2024)). Connector modules translate encoded features into the LLM’s input space, evolving from simple MLP projections in LLaVA (Liu et al., 2023, 2024a) to more sophisticated mechanisms employing learnable query tokens (Li et al., 2023) and added cross-attention layers (Alayrac et al., 2022). Training typically follows a multi-stage approach: alignment pre-training to bridge encoders with LLMs using large-scale paired datasets, followed by multimodal instruction tuning (Liu et al., 2024a; Dai et al., 2023) to teach the models to better understand natural language instructions from users and solve tasks. The field continues to expand, incorporating 3D point clouds (Xu et al., 2024), video (Zhang et al., 2023), and any-to-any modal generation (Wu et al., 2024). Despite this progress, effective integration of continuous numerical data, particularly time series, remains a challenge due to the lack of high-quality paired time series and text datasets, and the complex, highly contextual nature of time series signals. TsLLM aims to address this gap by building on the foundations of MLLM research and adapting it to time series analysis.
@@
Inspired by recent successes in natural language processing (Zhao et al., 2023) and computer vision (Awais et al., 2025), research in time series analysis is increasingly focused on developing dedicated Time Series Foundation Models (TSFMs). These models are pretrained on large, diverse numerical time series corpora to learn universal temporal representations, and apply them to solving a task. The transformer architecture is a common backbone, with “patching” being a key adaptation for efficiency and local feature extraction, as demonstrated by PatchTST (Nie et al., 2023) for forecasting. Architectural variants include encoder-only models like Moirai (Liu et al., 2024b) for universal forecasting, and decoder-only models such as TimesFM (Das et al., 2024) pretrained on billions of data points for strong zero-shot forecasting.
While the aim of this body of work is to develop foundation models for time series, the vast majority of the current literature focuses solely on forecasting. There are a few notable exceptions, namely UniTS (Gao et al., 2024), which aims for broader multi-task and multi-domain generalization on numerical data through a unified token representation that encodes the task specifications, and MOMENT (Goswami et al., 2024), which trains a backbone model along with lightweight decoders for multiple tasks. Despite significant advances in learning from a broad set of numerical time series, these TSFMs lack the world knowledge and language processing capabilities that LLMs have to interpret rich, unstructured textual context, which is TsLLM’s goal.
@@
Distinct from efforts in dedicated TSFMs, applications of pretrained LLMs to time series analysis are rapidly evolving. The core objective is to leverage the rich contextual understanding of LLMs while adapting text-optimized architectures for continuous numerical data (Zhang et al., 2024). While early work demonstrated feasibility by prompting LLMs with text-based time series for zero-shot forecasting (Gruver et al., 2023), limitations in numerical precision and efficiency (Spathis and Kawsar, 2024) spurred the development of more sophisticated representation strategies. These include discretizing continuous series via quantization (Ansari et al., 2024; Zhang et al., 2025) or vocabulary expansion (Wang et al., 2025; Xie et al., 2024) (although at the cost of potential information loss (Chatzigeorgakidis et al., 2024)), and structuring the input through patching or multi-scale decomposition (Zhou et al., 2023; Jin et al., 2023; Chan et al., 2024; Kowsher et al., 2024). However, these methods often process time series and text as distinct, sequential blocks, focusing primarily on forecasting.
Recent work addresses input representation limitations by enforcing structural consistency within the LLM’s latent space to bridge modality gaps (Hu et al., 2025; Liu et al., 2025). Other approaches unlock reasoning potential through conversational and multi-agent systems for enhanced performance and interpretability (Xie et al., 2024; Lee et al., 2025). While these methods advance multimodal integration, they typically treat time series as a static input to be reasoned about, generating textual summaries or predictions as the final output. TsLLM addresses these critical limitations by interleaving time series and text within the inputs and, crucially, the outputs of a pretrained LLM. This enables deep connections between modalities and provides new flexibility in how problems can be constructed and solved.
@@ 3 Methodology
We propose TsLLM, a multimodal generative model that augments a pretrained decoder-only LLM with time series perception and generation abilities. Our design goals are: (i) to represent time series with high information density while preserving temporal structure and numeric fidelity; (ii) to enable interleaved text-time series inputs and outputs in a single autoregressive model; and (iii) to train the resulting model with a curriculum that first aligns modalities and then teaches a broad set of time series tasks under natural language instructions. TsLLM consists of four main components: (1) a patch-based time series encoder-decoder, (2) a pretrained LLM, (3) a pair of fusion layers that inject time series embeddings into the LLM token stream and maps LLM hidden states back to time series patches, and (4) a multi-stage training pipeline, as illustrated in Figure 1.
Our unified framework reformulates all time series tasks, including forecasting, classification, anomaly detection, and general question-answering, as autoregressive generation problems over interleaved sequences of text and time series tokens. This approach enables few-shot generalization to new domains through natural language instructions and in-context examples, while supporting rich contextual information such as domain knowledge, metadata, and even structured tabular data. The following sections detail each component that are primary contributions of this work, namely, our patch-based time series encoding mechanism (Sec. 3.1), the architectural modifications for multimodal integration (Sec. 3.2), and our multi-stage training strategy (Sec. 3.3).
@@ 3.1 Time Series Encoding and Decoding
TsLLM depends on high-quality representations of time series signals for strong performance. However, current LLMs split numbers into multiple small tokens, making it hard to capture numerical relationships and resulting in inefficient, lengthy text-based sequences, with each number using four to eight tokens including delimiters (Spathis and Kawsar, 2023). This limits the amount of contextual information, e.g., covariates and few-shot examples, that can fit in the context window, and requires an alternative data encoding.
Patch-Based Encoding Strategy
We solve this representation challenge through a patch-based encoding mechanism that preserves temporal structure while achieving dramatic compression. Our approach segments time series into fixed-length patches of $p$ consecutive points, transforming each patch into a single dense representation. This design decision balances three competing objectives: (1) capturing sufficient local temporal patterns within each patch, (2) maintaining computational efficiency through reasonable sequence lengths, and (3) preserving fine-grained information necessary for precise analysis. The encoding process operates in two stages. First, each patch $\mathbf{x}_{i:i+p}$ undergoes local feature extraction through a learned MLP: $\mathbf{z}_{i}^{\text{local}}=\text{MLP}_{\text{patch}}(\mathbf{x}_{i:i+p})$, where the MLP captures patterns like trends, oscillations, and anomalies within the patch window. Second, these local representations are contextualized through a decoder-only transformer that models dependencies across patches: $\mathbf{z}_{i}=\text{TransformerEncoder}(\mathbf{z}_{1}^{\text{local}},...,\mathbf{z}_{n}^{\text{local}})_{i}$. The causal attention mechanism ensures compatibility with autoregressive generation, enabling the model to generate time series continuations naturally.
Architecture Details
Our encoder-decoder architecture employs several design choices optimized for time series representation. We use learned positional encodings rather than sinusoidal variants, allowing the model to adapt to different sampling frequencies and temporal scales. Pre-normalization and GELU activations improve training stability when handling diverse signal magnitudes. The patch size $p=32$ was selected through extensive ablation studies, providing an optimal trade-off: smaller patches preserve more detail but increase sequence length, while larger patches risk losing important local features.
For multivariate time series, we adopt a channel-independent encoding strategy where each univariate component is processed separately. This design choice, while seemingly restrictive, offers significant advantages: it naturally handles varying numbers of channels, enables transfer learning across different sensor configurations, and simplifies the alignment between time series and text modalities. The LLM backbone subsequently learns cross-channel relationships through its attention mechanism, leveraging its pre-trained capabilities for modeling complex dependencies.
@@
Real-world time series exhibit enormous variability in scale – from microvolts in neural recordings to millions in financial transactions. To ensure stable training across this range, we apply sequence-level z-score normalization before encoding: $\hat{\mathbf{x}}=(\mathbf{x}-\mu_{\mathbf{x}})/\sigma_{\mathbf{x}}$, where $\mu_{\mathbf{x}}$ and $\sigma_{\mathbf{x}}$ are computed over the entire input sequence. We pretrain the encoder-decoder pair using a $\beta$-VAE objective that balances reconstruction fidelity with representation quality. The loss function combines two terms:
$\mathcal{L}_{\text{VAE}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]-\beta\cdot\text{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))$
@@
where the reconstruction term (MSE loss) ensures accurate signal recovery, while the KL divergence term regularizes the latent space to be smooth and well-structured. We employ a sigmoid-based schedule for $\beta$, starting near zero to prioritize reconstruction accuracy, then gradually increasing to encourage disentangled representations. This scheduling strategy prevents the common “posterior collapse” problem in VAE training while ensuring high-quality reconstructions. The encoder-decoder is pretrained on 50 million sequences from the GIFT-Eval Pretrain dataset (Aksu et al., 2024), comprising 230 billion time series points across 88 diverse domains, including healthcare, finance, and industrial sensors. This large-scale pretraining ensures our representations generalize across domains, temporal resolutions, noise levels, and time horizons.

### v2

# TsLLM: Augmenting LLMs for General Time Series Understanding and Prediction

@@ Abstract
Time series data is fundamental to decision-making across many domains including healthcare, finance, power systems, and logistics. However, analyzing this data correctly often requires incorporating unstructured contextual information, answering domain-specific questions, and generating natural language explanations – capabilities that traditional time series models lack. While Large Language Models (LLMs) excel at contextual reasoning and knowledge integration, they struggle with numerical time series due to inefficient text-based representations and limited exposure to numerical data during pretraining. We address this gap by augmenting an LLM with specialized time series perception through a patch-based encoder-decoder architecture. We train this Time Series-augmented LLM (TsLLM) on a large corpus of over 25 billion tokens of interleaved time series and text spanning diverse tasks: forecasting with contextual information, question-answering, anomaly detection, classification, report generation, and more, all unified as next token prediction. This training enables TsLLM to leverage both its language understanding and newly acquired temporal reasoning capabilities. While not designed to surpass specialized models on traditional benchmarks, TsLLM demonstrates strong performance on tasks requiring the integration of time series analysis with natural language – capabilities that existing approaches cannot provide. It also exhibits strong zero-shot and few-shot performance, showing it can adapt to new data without additional training.
Keywords Time Series $\cdot$ LLMs $\cdot$ Multimodal $\cdot$ Temporal Reasoning
@@ 1 Introduction
Traditional time series models are effective for narrowly defined numerical tasks, mapping structured inputs to structured outputs; however, many real-world applications of time series analysis require more than numerical prediction – context is often crucial. Clinicians use electrocardiograms for diagnosis in the context of symptoms and lab results, analysts forecast while conditioning on news, and power grid operators balance electricity capacity and demand using load and frequency traces alongside weather forecasts, outage reports, and event calendars. Across healthcare, finance, scientific research, and many other fields, domain experts need systems that can understand both numerical time series data and all relevant contextual information, draw on domain knowledge, make predictions, and effectively communicate findings.
@@
Large language models (LLMs) possess the capabilities needed for these integrated tasks: contextual reasoning, few-shot learning from examples, domain knowledge integration, and natural language interaction. As a result, utilizing LLMs to perform certain time series analysis tasks holds considerable promise [1, 2]. However, their weak abilities to perceive numerical time series data prevent LLMs from effectively applying their capabilities to temporal analysis [3, 4]. Two structural issues explain this gap. First, text tokenization represents each numeric value with multiple tokens, inflating sequence length and obscuring temporal structure; this limits the amount of signal and context that fits in context windows and hinders pattern recognition. A single 10 second electrocardiogram used in our experiments (12 leads, 100Hz) requires 84,000 tokens using a standard tokenizer – longer than the maximum context length of many models. Second, our analysis of LLM pretraining corpora reveals they contain little high-quality time series data, so LLMs lack the exposure needed for temporal analysis.
@@
To address these limitations, we introduce TsLLM, a large language model augmented with dedicated time series perception that enables unified, interleaved text-time series reasoning and generation. TsLLM adds a high-fidelity, patch-based encoder-decoder that is pretrained with a $\beta$-VAE objective to produce compact, information-dense representations of temporal signals, along with a scale-aware encoding that decouples shape from magnitude so the model can reason about absolute scale while maintaining numerical stability. A lightweight cross attention-based adapter then maps these time series representations into the LLM’s embedding space to bridge the domain gap. TsLLM’s inputs and outputs can interleave text and time series segments, and the model treats all tasks as next-token prediction, unifying these modalities. This preserves the pretrained LLM’s language ability while adding precise temporal perception. TsLLM’s unified formulation of language and time series modeling enables functions that traditional time series models and text-only LLMs generally do not provide: contextual forecasting conditioned on unstructured information, question answering over numerical time series data, and natural language guided time series synthesis, as well as standard time series tasks: forecasting, classification, anomaly detection and more, without using task-specific heads. Few-shot in-context prompting with interleaved examples further allows the model to adapt to new datasets and tasks without additional fine-tuning.
@@
Empirically, TsLLM demonstrates strong performance on tasks that require integrating temporal signals with language. TsLLM significantly outperforms frontier LLMs on time series question-answering (QA) benchmarks, matches or exceeds specialized systems for QA on electrocardiogram data, and beats both LLM-based methods and traditional models on a contextual forecasting benchmark. It does this while maintaining its performance on standard text-only LLM evaluations. On traditional time series benchmarks, TsLLM is competitive with dedicated models on forecasting and classification, even though it is not optimized for purely numerical tasks. Ablations show that few-shot prompting provides significant gains in most settings, and that the performance of TsLLM scales with the size of the pretrained LLM it is built on, suggesting that its performance can be improved by scaling up further. The primary contributions of this work are:
@@
    A large corpus of multimodal interleaved text and numerical time series sequences, comprising over 25 billion tokens in total, which were collected, augmented, and synthesized for training TsLLM.
@@
    We will release both the data and code for TsLLM to enable further research in this area.
@@ 2 Related Works
Time Series Foundation Models (TSFMs) are pretrained on large, diverse numerical time series corpora to learn universal temporal representations [5, 6]. Most TSFMs focus primarily on forecasting, though notable exceptions exist. UniTS [7] aims for broader multi-task and multi-domain generalization through a unified token representation that encodes task specifications, while MOMENT [8] trains a shared backbone with lightweight task-specific decoders. However, all these models operate exclusively on numerical inputs and outputs, lacking the ability to incorporate unstructured textual context or generate natural language explanations. TsLLM is designed to complement rather than compete with TSFMs: it targets problems requiring integrated temporal and linguistic reasoning, while TSFMs remain well-suited for pure numerical tasks.
LLM-Based Time Series Approaches
Given the limitations of both text-only LLMs and purely numerical TSFMs, a growing body of work attempts to bridge these modalities by applying LLMs to time series analysis [2, 1]. Early approaches worked within these constraints by prompting LLMs with numerical sequences as strings [9, 10], but suffered from precision loss and inefficiency. Evidence suggests that LLMs fundamentally struggle with time series understanding and that architectural augmentation, rather than prompting alone, is necessary [4, 11].
Subsequent work aimed to improve performance using better time series encoding and alignment with the LLM, including reprogramming approaches like Time-LLM [12], which align time series with text prototypes while keeping the LLM frozen, and others [13, 14]. These methods are able to achieve strong performance on traditional time series tasks, but do not leverage the LLM’s language generation capabilities and require dataset-specific training, limiting their value.
More recent work has explored methods to better utilize the text understanding and generation capabilities of LLMs. Some convert time series to plots and leverage vision-language models [15, 16], which avoids the need for specialized time series encodings, but sacrifices precision, and these methods do not train the LLM. ChatTime [17] quantizes continuous values into discrete tokens, and fine-tunes an LLM using these tokens for forecasting and question-answering. ChatTS [18] uses an MLP encoder to map time series patches directly into the LLM’s embedding space, and trains on synthetic question-answering data.
However, these methods all have notable shortcomings. Due to the absence of large, standardized datasets for mixed time series and text tasks, these methods train on relatively small, often entirely synthetic datasets. This limits their ability to generalize to other datasets and tasks effectively. Many use quantized or image-based time series encodings, which limit their perception abilities and precision. Some can utilize time series inputs, but cannot generate time series outputs, limiting the tasks they can perform. TsLLM aims to address the limitations of existing methods by using a high-fidelity encoder-decoder for time series inputs that is pretrained for good performance, supporting arbitrary interleaved sequences of time series and text to enable it to solve all tasks involving these modalities, and training on a new massive collection of these mixed sequences totaling over 25 billion tokens.
@@
An adjacent line of research is in using LLM-based models for time series tasks [2]. Some of these methods are also TSFMs, but many are not. The core motivation of this approach is to use pretrained LLMs, which are extremely capable sequential prediction models as a foundation for time series prediction. While early work demonstrated feasibility by prompting LLMs with text-based time series for zero-shot forecasting [9], limitations in numerical precision and efficiency [19] spurred the development of more sophisticated representation strategies. These include discretizing continuous series via quantization [20, 21] or vocabulary expansion [17, 18], and structuring the input through patching or multi-scale decomposition [22, 12, 13, 14]. However, these methods generally do not use the text understanding and generation capabilities of LLMs at all, or do so in very limited ways, negating their advantages. They also focus primarily on forecasting, and often do not train the LLM itself, only encoder and decoder layers. Some more recent work has aimed to better utilize the text understanding and generation capabilities of LLMs [18, 23], but use weaker text-based time series representations and are more limited in scope.
@@ 3 Methodology
We propose TsLLM, a multimodal generative model that augments a pretrained decoder-only LLM with time series perception and generation abilities. Our design goals are: (1) build good representations of time series signals that are structured and compact without significant information loss; (2) allow for a mixture of text and time series in inputs and outputs to provide flexibility in constructing tasks; and (3) ensure the model has sufficient training to allow it to perform well across a highly diverse set of tasks and data. To meet these goals, TsLLM consists of four main components: (1) a patch-based time series encoder-decoder, (2) a pretrained LLM, (3) a pair of fusion layers that inject time series embeddings into the LLM token stream and maps LLM hidden states back to time series patches, and (4) a multi-stage training pipeline. These components are illustrated in Figure 1.
Our unified framework reformulates all time series tasks, including forecasting, classification, anomaly detection, and general question-answering, as autoregressive generation problems over interleaved sequences of text and time series tokens. This approach enables few-shot generalization to new domains through natural language instructions and in-context examples, while supporting rich contextual information such as domain knowledge, metadata, and even structured tabular data. The following sections detail each component that are primary contributions of this work, namely, our time series encoding mechanism (Sec. 3.1), the architectural modifications to the LLM (Sec. 3.2), and our training strategy (Sec. 3.3).
@@ 3.1 Time Series Encoding and Decoding
TsLLM depends on high-quality representations of time series signals for strong performance. However, current LLMs split individual numbers into multiple tokens, making it hard for models to capture numerical relationships and resulting in extremely long, inefficient inputs, with each time point using four to eight tokens (including delimiters) [24]. This limits the amount of time series and contextual information, particularly covariates and few-shot examples, that can fit in the context window, and makes training very computationally expensive. An alternative time series encoding is needed.
We solve this representation challenge through a patch-based encoding mechanism that preserves information while achieving dramatic compression. Our approach segments time series into fixed-length patches of $p$ consecutive points, and transforms each patch $\mathbf{x}_{i:i+p}$ into a single dense representation through a learned MLP: $\mathbf{z}_{i}^{\text{local}}=\text{MLP}_{\text{patch}}(\mathbf{x}_{i:i+p})$. The patch size $p=32$ was selected through extensive ablation studies, providing an optimal trade-off: smaller patches preserve more detail but increase sequence length, while larger patches risk losing important local features. Each component in a multivariate signal is encoded independently, allowing for flexibility in the number of components. The resulting univariate signals are fed into the LLM sequentially, allowing it to learn the relationships itself.
@@
Real-world time series exhibit enormous variability in scale – from microvolts in neural recordings to millions in financial transactions. To ensure stable training across this range, we apply sequence-level z-score normalization before encoding: $\hat{\mathbf{x}}=(\mathbf{x}-\mu_{\mathbf{x}})/\sigma_{\mathbf{x}}$, where $\mu_{\mathbf{x}}$ and $\sigma_{\mathbf{x}}$ are computed over the entire input sequence.
@@
We pretrain the encoder-decoder pair using a $\beta$-VAE objective that balances reconstruction fidelity with representation quality. The loss function combines two terms: $\mathcal{L}_{\text{VAE}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]-\beta\cdot\text{KL}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))$ where the reconstruction term (MSE loss) ensures little information is lost, while the KL divergence term regularizes the latent space to be smooth and well-structured. We employ a sigmoid-based schedule for $\beta$, starting near zero to prioritize reconstruction accuracy, then gradually increasing to encourage disentangled representations. This scheduling strategy reduces the common posterior collapse problem in VAE training while ensuring high-quality reconstructions. The encoder-decoder is pretrained on 50 million sequences from the GIFT-Eval Pretrain dataset [25], comprising 230 billion time series points across 88 diverse domains, including healthcare, finance, and industrial sensors. This large-scale pretraining ensures our representations generalize across domains, temporal resolutions, noise levels, and time horizons.
Overall, this design successfully balances three competing objectives: (1) capturing sufficient local temporal patterns within each patch, (2) maintaining computational efficiency through reasonable sequence lengths, and (3) preserving fine-grained information necessary for precise analysis.